# Aristotle’s Second-Order Logic

### v11

@@ Abstract
This paper has two aims. The first aim is to argue that an extension of second-order logic in its natural deduction presentation is the most adequate framework to formalize the metalogic of the Analytics and the logic implicit in the Topics as well as the reasoning present in other works of Aristotle. The second aim is to provide evidence for the thesis (defended by Bobzien[7]) that ancient logic had sufficient means to deal effectively with multiple generality.
@@ Introduction
In the last decades there has been an increased interest in the interpretation of Aristotle’s Organon using the framework of modern mathematical logic, the pioneers of this approach being found, for instance, among the Polish school of logicians and philosophers of the first half of the twentieth century. Similar research has also been carried out for other schools of ancient logic, notably for Stoic logic and ancient Indian logic. In this context some researchers have argued that modern and contemporary formal and philosophical approaches to logic were already present among ancient logicians. Thus we go beyond merely using modern formal structures and methods as interpretative tools but seek evidence (as far as the unfortunately fragmentary state of the material is concerned) that concepts and structures very similar to these were already present therein [^2] Some arguments for the legitimacy of deploying modern formal concepts to interpret historical philosophical systems (what is called the hermeneutical ’virtuous circle’) are found in [13]. We add here that the ability of an ancient philosophical text to be translated in a more-or-less direct way into a variant of a standard modern formal system is in itself remarkable. From the start the choice of the most adequate system is not obvious and the situation is usually one of great (even unsurmountable) difficulties.. For example in the Prior Analytics Aristotle seems to make the distinction between a term and its extension, between proof theoretic arguments and and model theoretic concepts such as soundness by making use of the construction of models to serve as counterexamples. The modal syllogistic can be axiomatized and given a possible world semantics[17]. Stoic logic can be interpreted as a sequent calculus for propositional logic - in fact as a substructural relevance logic[6].
The present work’s goal is to look at the metalogic (the reasoning in the arguments Aristotle gives, for example, about the conversion of propositions or the validity and non-validity of syllogisms) employed in the Analytics and the logic implicit in the Topics and to argue that: i) it corresponds to an extension of second-order logic - presented in Jaśkowski333in this paper when we refer to natural deduction we assume that the Jaśkowski style presentation is meant. Contrary to Gentzen’s sequent-style presentation this version resembles more closely how actual reasoning is carried out. We also deploy its linearized version (also called Lemmon style). natural deduction style - and many of Aristotle’s statements and proofs involve nested quantifiers and non-monadic predicates, ii) that Aristotle was conscious of the natural deduction rules and precise in their application, iii) to give a first sketch of a formal reconstruction and interpretation of the system of argumentation (in which definitions or other kinds of hypothesis were tested and possibly refuted) for which the Topics was originally written.
The paper is organized as follows. In section 2 we focus on the metalogic of the Analytics and show that its most natural interpretation is as an extension of second-order logic which we call Aristotelian second-order logic (ASOL). ASOL is second-order logic extended with a collection of second-order $n$-ary predicates having as arguments predicate terms and possibly individual terms. We also include some remarks on how the formalization of the modal syllostic in [17] can be embedded into ASOL. In section 3 we argue that Aristotle consciously employs natural deduction rules for universal and existential quantification. In the Analytics the key second-order binary predicates are posited to be definable (at least in the non-modal setting) in terms of strict second order logic, which expresses Aristotle’s proto-Fregean extensionalism. In section 4 we study some key aspects of Aristotle’s Topics and show that ASOL is the best choice for representing the underlying logic. Here we require a richer collection of second-order predicates (expressing the predicabilia) as compared to the Analytics; we do not have any ’extensionalist’ reduction of the main second-order predicates to pure second-order logic, more specifically, to relations between first-order predication. A topic is a sentence ASOL in general universally quantified over most of the monadic predicate variables occurring in the sentence. These universally quantified monadic predicate variables are then supposed to be instantiated according to the dialectical needs at hand. Although the structure of the Topics consists mainly in a rather tedious enumeration of topics, the most purely logical among them can be organized into groups wherein all the topics in a group are logical consequences of a certain paradigmatic topic. Often Aristotle indicates explicitly that for two closely related topics (mentioned in succession) one is the consequence of the other. This circumstance provides evidence to gauge the strength and nature of the logic implicitly present. In section 5 we look at book V which deals with property (idion) and discuss some of the challenges facing a formalization of the topics included therein as well as analyzing Aristotle’s treatment of superlatives in book VII, 152a.
In the concluding section we touch upon the significance of our approach for the history of the development of logic and in particular the connection between this interpretation of the Organon and Frege’s logic. We end with a discussion about how the system of debate for which the Topics was written might be reconstructed and interpreted according to ASOL.
In order to provide more context for section 4 in appendix 1 we list a formalization of an important part of the topics of book IV, focused on the relation between genus and species. In appendix 2 we illustrate the dialectical process for which the Topics was written by exhibiting a ASOL formalization of a key section of Plato’s Lysis. Appendix 3 presents a minimal set of axioms for a formalization of the first proposition of Euclid and appendix 4 includes the definition of ASOL and its system of natural deduction. This paper does not assume any technical knowledge of modern formal logic beyond an understanding of the language and rules of (extended) second-order logic presented in appendix 4.
@@
In this paper we will work in Aristotelian Second-order logic (ASOL) which is standard second-order logic extended with a collection of second-order $n$-ary predicates having as arguments predicate terms and possibly individual terms. We consider its standard natural deduction presentation. See appendix 4 for details. In this paper we consider only classical rather than intuitionaistic ASOL, that is, we have the rule $\bot_{c}$.
@@
There have been many attempts to formalize Aristotle’s syllogistic expounded in the Prior Analytics in both its strictly assertoric and modal variants (see [17] for references). Some approaches involve monadic (modal or otherwise) first-order logic and other approaches use variants (modal of otherwise) of propositional logic. Additionally many authors assume that Aristotle reasons much like a modern logician by constructing models to refute the validity of certain syllogistic figures. The above approaches usually set up a minimal set of axioms together with a deductive system and attempts to show the agreement (at in least for most cases) between the derivable syllogisms and those considered valid by Aristotle (how Aristotle himself derived or justifies the validity or non-validity playing a secondary role). Our concern is different. We are interested in the structure of the proofs in the Analytics themselves, how Aristotle argues for the validity or non-validity of a certain figure or propositional conversion. The modern division between metalogic and object logic does not seem to be, at first glance, so evident here. Aristotle does not simply posit some axioms and a deductive system and show how some figures are valid and some are not (with the use of models). Rather it appears that what we now call metalogic and object logic form an inextricable whole. So the best formalization must be directly about this whole, about the logical and argumentative structure of the Analytics themselves. A noteworthy characteristic of the Analytics is the presence of variable letters ranging over Aristotelian terms which (so we argue) usually correspond to monadic predicate terms in second-order logic. There is a distinguished class of entities for which it is meaningful to ask if such entities fall or do not fall under such monadic predicate term - these we call individuals. Thus we can ask of a monadic predicate term is predicated or not of a given individual.
@@
It is patent that a formalization of Aristotle’s reasoning must be a term-logic, that is, terms must be the basic building blocks - the most basic division being between individual terms (corresponding to ’first substances’) and non-individual terms. Sentences (for the moment we consider here only ’atomic’ sentences (heis logos, De Int. 17a16-17) are then built up from combinations of terms in a specified manner. Predication - and in Aristotle there are many distinct notions of predication hiding under the verb esti - corresponds to combinations of terms tagged according to the kind of predication being considered. Thus if ① corresponded to predicating a non-individual term $A$ of an individual term $a$ (let us say accordingo to ’essence’) and ② corresponded to predicating a non-individual term $B$ of another $A$ according to ’essence’ then grammatically (for SOV languages like Ancient Greek) and notationally the most natural representation would be aA① and AB②. Such a representation is necessarily if we wish to formalize correctly’s Aristotle’s essentially intensional logic: the predication of definition and the predication of property (idion) both imply that the terms have equal ’extension’ (in the sense of individuals falling under them) but the types of predication are different as are the terms: ’being able to laugh’ is not the same term as ’rational animal’ but both are co-extensional with ’man’. Thus we have grounds for holding that an extension of second-order logic like ASOL (in which the circled number tags correspond to our second-order predicates and ① to simple concatenation in reverse order $Aa$) is the most adequate minimal formalization for the logical-argumentative apparatus in the Organon and even, as we shall discuss later, in the Physics. What we wrote above applies to ’atomic’ sentences. But do not modern formal systems with their full inductive definitions of formulas have greater complexity and precision than can be reconstructed from Aristotle’s system embodying an only seemingly rudimentary account of complex sentences ? Against this worry we reply that the fact of the matter is that rigorous structural inductive definitions of sentences are not an exclusive possession of modern logic. A perfectly ’modern’ definition of the formulas of what we would now consider an extension of modal propositional logic (a multi-modal propositional logic) is found in Ockham’s Summa Logicae[14][79-80]. Going back to antiquity there is strong evidence for such inductive definitions in Stoic Logic[6].
@@
The four basic kinds of protasis dealt with in the Analytics can be seen as corresponding to four distinct second-order binary predicates, even if we grant that such second-order predicates can be defined without such second-order predicates. Syllogisms correspond very naturally to complex sentences in ASOL of a certain structure:
$\forall A,B,C(\phi(A,B)\&\psi(B,C)\rightarrow\rho(A,C))$
@@
where $\phi,\psi$ and $\rho$ range among certain atomic second-order binary predicates (the order of the free variables is not necessarily the order of the free variables in the atomic binary predicate). For example, the Barbara syllogism corresponds to:
@@
where the second-order atomic binary predicate $\kappa XY$ is to be read as ’all $X$ is $Y$’.
From a formal point of view this is exactly what Aristotle is expressing and what Aristotle makes use of in the structure of the proofs in the Analytics. Aristotle certainly seems to assume the following reductive axiom (or rather definition) in the proofs:
@@
Thus $\kappa$ is extensional. But Axiom (A) has to be tweaked as it is unlikely that Aristotle would have accepted $\kappa AA$.
We introduce an extensional second-order equality predicate
@@
It is well known, Aristotle seems to assume that all terms have non-empty extension (even elements of the such extensions are classed as ’non-being’). Thanks to ASOL we can state this directly as an axiom
$\forall A\exists xAx$
Another fundamental second-order binary predicate is $\mu XY$ which expresses ’some $X$ is $Y$’. We have the following axiom or definition
@@
which states that $\mu$, like $\kappa$, is extensional. Again we mention the problem that we have to exclude $\mu AA$ and this leads us to
@@
We can follow the approach of [17] and introduce a second-order monadic function to represent term-negation (de re negation) which we denote by $A^{c}$. But thanks to our ASOL framework we can now define it directly
@@
Note that from (NEmp) we deduce that $\forall A\exists x\sim Ax$. Finally we can define $\hat{\kappa}XY\leftrightarrow\kappa XY^{c}$ and $\hat{\mu}XY\leftrightarrow\mu XY^{c}$. Alternatively we could introduce $\hat{\kappa}$ and $\hat{\mu}$ directly with the expected axioms, for instance
@@
We now present some evidence that the ASOL approach captures the structure of the metalogical proofs in the Prior Analytics.
Aristotle in 25a announces the conversion rule
$\forall A,B(\hat{\kappa}AB\rightarrow\hat{\kappa}BA)$
Let us pay close attention to how he proves it in the text. Aristotle’s argument is closely paraphrased as follows:
Let $A$ and $B$ be terms and assume that $\hat{\kappa}AB$. Assume per absurdum that $\sim\hat{\kappa}BA$. Then by definition of $\hat{\kappa}$, $\sim\forall x(Bx\rightarrow\sim Ax))$, that is
$\exists x\sim(Bx\rightarrow\sim Ax)$
were we note the implied knowledge of propositional logic. Using again this knowledge the above is considered equivalent to
$\exists x(Bx\&Ax)$
Now we argue that Aristotle is actually deploying the $\exists^{1}$ rule, so he considers a witness $C$ (we use $y$) and introduces the hypothesis $By\&Ay$. But by hypothesis $\hat{\kappa}AB$, that is $\forall x(Ax\rightarrow\sim Bx)$. But then using $\forall^{1}$ with $y$ we get
$Ay\rightarrow\sim By$
[…]
Aristotle’s goal in chapters 4 - 7 of the Prior Analytics is to find all valid sentences of the form (Syll). While the logically valid syllogisms can be seen as being deduced in second-order logic from the axioms (mainly (Econ)), invalid sentences are eliminated by exhibiting counterexamples. Metalogically we must be able to express that a certain sentence is not a syllogism, for example in 26a:
$\sim\forall A,B,C(\kappa AB\&\hat{\kappa}CB\rightarrow\kappa CA)$
which Aristotle takes to be equivalent to
$\exists A,B,C\sim(\kappa AB\&\hat{\kappa}CB\rightarrow\kappa CA)$
that is
$\exists A,B,C(\kappa AB\&\hat{\kappa}CB\&\sim\kappa CA)$
This in turn is shown by producing concrete witnesses. In this case $A=animal,B=man,C=stone$. Aristotle assumes as valid
$\kappa\,man\,animal\&\hat{\kappa}\,stone\,man\&\sim\kappa\,stone\,animal$
and his reasoning can be interpreted as applying repeatedly second-order $\exists^{2}I$ to obtain (b) which is then seen to be logically equivalent to (a). What we have seen exhibits non-trivial reasoning with quantifiers over non-monadic (second-order) predicate. A nice aspect of the ASOL intepretation is that we do not have to leave ASOL to formalize Aristotle’s counterexamples.
Thus second-order logic with natural deductive rules appear to be a suitable framework to interpret the metalogic, the argumentation in the Prior Analytics themselves.
There are other passages where Aristotle can be seen as using a mixture of universal and existential quantification over monadic predicate variables. In the Posterior Analytics Book I, chapter 15, Aristotle gives the definition of immediate predication which can be expressed directly in ASOL as
$\forall A,B(\hat{\kappa}^{\ast}AB\leftrightarrow\kappa AB\&\sim\exists C(\kappa AC\&\kappa CB))$
We now consider briefly how the modal syllogistic could be treated in the framework of ASOL. To this end we consider an additional second-order monadic function $\square X$. This takes a monadic predicate term $X$ and yields a monadic predicate term $\square X$. We posit the following axioms:

### v17

@@ Abstract
This paper has two goals. The first goal is to show how an extension of second-order logic is a natural framework to formalize portions of Aristotle’s Topics and to bring to the foreground the logical, linguistic and philosophical interest of this work. Aristotelian logic and its related traditions in antiquity are often held to have been equivalent to monadic predicate logic and as such inadequate to formalize mathematics as well as scientific and philosophical discourse in general. The second goal of this paper is to argue that on the contrary the logical theories of Aristotle (which we argue correspond to a variant of natural deduction) and ancient authors such as Galen and Boethius were in fact quite sufficient to account for the logically complex expressions and reasoning involving multiple generality fundamental to the aforementioned disciplines.
@@ Introduction
This paper will be organized as follows. Sections 2-6 are dedicated to the formalization and discussion of Aristotle’Topics using an extension of second-order logic. In sections 7-9 we argue that the logical theory of Aristotle and some ancient authors was fundamentally equivalent to a variant of natural deduction which was applied to formulas of any complexity. In section 2 we present the language and rules of an extension of classical second-order logical (ESOL) which will serve as the foundation for sections 3-6. In section 3 we discuss briefly some aspects of metalogic of the Analytics and the formal structure of ancient logic which will be followed by a fuller discussion in section 9. In section 4 we study some key aspects of Aristotle’s Topics and argue that ESOL is a good choice for representing the underlying logic. Here we require a richer collection of second-order predicates (expressing the predicabilia) as compared to the Analytics; we do not have a ’extensionalist’ reduction of the main second-order predicates to formulas involving only first-order predication. A topic is a sentence of ESOL in general universally quantified over most of the monadic predicate variables occurring in the sentence. These universally quantified monadic predicate variables are then supposed to be instantiated according to the dialectical needs at hand [^2] The definition of topic (locus) given by Boethius in De topicis differentiis agrees with this. The definition ’Locus autem sedes est argumenti, uel id unde ad propositam quaestionem conueniens trahitur argumentum’ suggests directly a second-order universal quantified sentence which can be instantiated to yield an inference scheme to be used according to the ’predicates’ at hand, the proposita questio.. Although the structure of the Topics consists mainly in a rather tedious enumeration of topics, the most purely logical among them can be organized into groups wherein all the topics in a group are logical consequences of a certain paradigmatic topic. Often Aristotle indicates explicitly that for two closely related topics (mentioned in succession) one is the consequence of the other. In section 5 we look at book V which deals with property (idion) and discuss some of the challenges facing a formalization of the topics included therein as well as analyzing Aristotle’s treatment of superlatives in book VII, 152a. In section 6 we end with a discussion about how the system of debate for which the Topics was written might be reconstructed and interpreted according to ESOL. In order to provide more context for sections 4-5, in appendix 1 we list a formalization of an important subset of the topics of book IV, focused on the relation between genus and species. In section 7 we set up a variant of ESOL in which all quantifiers are bounded by some formula and the elimination rules for $\exists$ are replaced by rules involving an indefinite version of the Peano operator. The resulting system is arguably closer to the logical and inferential mechanisms of natural language and provides a good foundation to argue for the main thesis of the second part of the paper. A common view is that whereas the deductive mechanisms and logical expressivity of ancient Greek mathematics corresponds to the full power of modern first or higher order predicate logic, it was nonetheless the case that Aristotle’s theory of deduction (identified with the theory of the syllogism of the Prior Analytics ) captures only monadic predicate logic and was thus radically inadequate for the logical structure and deductive mechanism of mathematics as well as that of scientific and philosophical discourse in general. In sections 8 and 9 we challenge the above view and argue that the logical theories of Aristotle and some authors of late antiquity (Galen and Boethius) were in fact sufficient to account for the logically complex expressions and deductive processes of ancient mathematics, science and philosophy. To this end in section 8 we show first that Aristotle not only makes use of the four quantifier rules for ESOLD adequate for reasoning about quantified sentences but does so in a philosophical reflected way, that is, the quantifier rules are part of his theory of logic. We also discuss his use of reductio ad absurdum and the rules related to implication and inference. We then show that these rules are deployed by Aristotle for sentences containing multiple quantifiers over relations and furthermore present evidence for such a deployment in Galen. Section 9 is dedicated to discussing evidence for a formal theory of multiple generality in Boethius. For Aristotle we refer to the edition of the Greek text in [1] and [2] for the Topics, [19] for the Prior and Posterior Analytics and [20] for the Physics.
Extended Second-Order Logic
The language of extended second-order logic (ESOL) is as follows. We are given a collection of individual variables $x,y,z,$... and individual constants $a,b,c,$... together with a collection of $n$-ary predicate term variables $X^{(n)},Y^{(n)},Z^{(n)},$... (we omit the superscript if the context is clear) for $n\geq 0$. Furthermore we are given a collection of predicate term symbols $\alpha^{(\tau)},\beta^{(\tau)},\gamma^{(\tau)},$... where $\tau$ consists of a (possibly empty) list $(n_{1},...,n_{m})$ of integers $n_{i}\geq-1$ called a signature, where we likewise omit the superscript if the context is clear. It is also convenient to consider that we have a collection of function symbols $f^{(\tau,n)},g^{(\tau,n)},$... where $\tau$ is a signature and $n\geq-1$. For n=1n=-1 these are called first-order functional symbols. Otherwise they are called second-order function symbols.
If $\alpha^{(\tau)}$ is a predicate term symbol with $\tau=(n_{1},...,n_{m})$ non-empty then $\alpha\mathfrak{t}_{1}....\mathfrak{t}_{m}$ is a formula where if $n_{i}\geq 0$ then $\mathfrak{t}_{i}$ is a $n_{i}$-predicate term and if $n_{i}=-1$ then $\mathfrak{t}_{i}$ is an individual term. If $\tau$ is empty then $\alpha$ is considered a 0-predicate term and we write the corresponding formula as $\alpha()$.
If $\Phi$ is a formula and $x_{1},...,x_{n}$ is a list of individual variables (possible empty) then $\lambda x_{1}...x_{n}\Phi$ is a $n$-predicate term (for the empty list we put $n=0$ and write $\lambda\Phi$).
If $X^{(n)}$ is a predicate term variable then it is a $n$-predicate term. If $\alpha^{(\tau)}$ is a predicate term symbol such that $\tau=(k_{1},...,k_{n})$ with $k_{i}=-1$ then $\alpha^{(\tau)}$ is a $n$-predicate term333we can also call $\alpha$ in this case a $n$-predicate constant.
An individual constant or variable is an individual term.
If $f^{(\tau,n)}$ is a function symbol then $f\mathfrak{t}_{1}...\mathfrak{t}_{m}$ is an individual term (for n=1n=-1) or $n$-predicate term (for $n\geq 0$) where $\mathfrak{t}_{1}...\mathfrak{t}_{m}$ satisfy the condition relative to $\tau$ as above.
The standard natural deduction rules consist of the standard rules for the first-order classical predicate calculus:
The rules are as follows:
$\Phi$$\Psi$  &I\&I $\Phi\,\&\,\Psi$ $\Phi\,\&\,\Psi$  $E_{R}$   $\Phi$ $\Phi\,\&\,\Psi$  $E_{L}$   $\Psi$
   $\Phi$  $\vee I_{R}$ $\Phi\,\vee\,\Psi$    $\Psi$  $\vee I_{L}$ $\Phi\,\vee\,\Psi$ $\Phi\lor\Psi$    $\Phi$] $\Upsilon$     $\Psi$] $\Upsilon$  $\vee E$      $\Upsilon$
[…]
$\bot$  $\bot_{i}$ $\Phi$ $(\sim\Phi)$   $\bot$    $\bot_{c}$   $\Phi$
the four first-order quantifier rules:
  $\Phi$  $\forall^{1}I$ $\forall x\Phi^{y}_{x}$ $\forall x\Phi$  $\forall^{1}E$ $\Phi^{x}_{t}$
$\Phi^{x}_{t}$  $\exists^{1}I$ $\exists x\Phi$ $\exists x\Phi$    $(\Phi^{x}_{y})$   $\Psi$  $\exists^{1}E$     $\Psi$
We have the analogous second-order quantifier rules
   $\Phi$  $\forall^{2}I$ $\forall X^{n}\Phi^{Y^{n}}_{X^{n}}$ $\forall X^{n}\Phi$  $\forall^{2}E$   $\Phi^{X^{n}}_{T^{n}}$
  $\Phi^{X^{n}}_{T^{n}}$  $\exists^{2}I$ $\exists X^{n}\Phi$ $\exists X^{n}\Phi$    $(\Phi^{X^{n}}_{T^{n}})$    $\Psi$  $\exists^{2}E$    $\Psi$
with analogous provisos.
Finally there are introduction and elimination rule for $\lambda$-predicate terms for which the reader may consult [14]. In Prawitz[14] versions 1 and 2 are proof-theoretically equivalent (for formulas without $\lambda$), the difference being that in version 1 the rules $\lambda I$ and $\lambda E$ are incorporated into second-order quantifier rules [^5] Why do we consider this system as being an extension of second-order logic rather than a fragment of third-order logic ? We are of the view that what defines the ’order’ of a logic is what is quantified over and what kind of corresponding variable there is. First-order logic quantifies over individuals and has individual variables. Second-order logic quantifies over entities corresponding to $n$-ary predicates with individual arguments and has the corresponding $n$-ary predicate term variables. Third-order logic thus would have to quantify over entities corresponding to predicates which can involve both individual arguments and $n$-ary predicates of individuals arguments and furthermore it would require variables to represent these last. Such a quantification and such variables are lacking in this system although the predicate term symbols $\alpha^{(\tau)}$ can be considered in some sense as being ’third-order’ when involving $n$-predicate term arguments.. In this paper we only consider classical logic (i.e. we always have the rule $\bot_{c}$).
@@
3.1 Logic and language
@@
There have been many attempts to formalize Aristotle’s syllogistic expounded in the Prior Analytics in both its strictly assertoric and modal variants (see [16] for references). Some approaches involve monadic (modal or otherwise) first-order logic and other approaches use variants (modal of otherwise) of propositional logic. Additionally many authors assume that Aristotle reasons much like a modern logician by constructing models to refute the validity of certain syllogistic figures. The above approaches usually set up a minimal set of axioms together with a deductive system and attempt to show the agreement (at in least for most cases) between the derivable syllogisms and those considered valid by Aristotle (how Aristotle himself derived or justifies the validity or non-validity playing a secondary role). Our concern in this section is different. We are interested in the structure of the proofs in the Analytics themselves, how Aristotle argues for the validity or non-validity of a certain figure or propositional conversion - an approach reminiscent of the modern division between metalogic and object logic. A noteworthy characteristic of the Analytics is the presence of variable letters ranging over Aristotelian terms which (so we argue) usually correspond to monadic predicate terms in second-order logic. There is a distinguished class of entities for which it is meaningful to ask if such entities fall or do not fall under such monadic predicate terms - these we call individuals. Thus we can ask if a monadic predicate term is predicated or not of a given individual.
@@
It is patent that a formalization of Aristotle’s reasoning must be a term-logic, that is, terms must be the basic building blocks - the most basic division being between individual terms (corresponding to ’first substances’) and non-individual terms. Sentences (note that the heis logos, De Int. 17a16-17 cannot be identified with ’atomic’ sentences in the modern sense as they may contain connectives) are then built up from combinations of terms in a specified manner. Predication - and in Aristotle there are many distinct notions of predication hiding under the verb esti - corresponds to combinations of terms tagged according to the kind of predication being considered. Thus if ① corresponded to predicating a non-individual term $A$ of an individual term $a$ (let us say according to ’essence’) and ② corresponded to predicating a non-individual term $B$ of another $A$ according to ’essence’ then grammatically (for SOV languages like Ancient Greek) and notationally the most natural representation would be aA① and AB②. Such a representation is necessarily if we wish to formalize correctly Aristotle’s essentially intensional logic: the predication of definition and the predication of property (idion) both imply that the terms have equal ’extension’ (in the sense of individuals falling under them) but the types of predication are different as are the terms: ’being able to laugh’ is not the same term as ’rational animal’ but both are co-extensional with ’man’. Thus we have grounds for holding that an extension of second-order logic (in which ② corresponds to a predicate term symbol of signature $(1,1)$ and ① to simple concatenation in reverse order $Aa$) is an adequate minimal formalization for the logical-argumentative apparatus in the Organon. What we wrote above applies to ’atomic’ sentences. But do not modern formal systems with their full inductive definitions of formulas have greater complexity and precision than can be reconstructed from Aristotle’s system embodying an only seemingly rudimentary account of complex sentences ? Against this worry we reply that the fact of the matter is that rigorous structural inductive definitions of sentences are not an exclusive possession of modern logic. There is strong evidence for such inductive definitions in Stoic Logic[5] and specially in Book I 1173D-1176D of Boethius’ De topicis differentiis which will be discussed in detail in section 9. Finally we mention that perfectly ’modern’ definition of the formulas of what we would now consider an extension of modal propositional logic (a multi-modal propositional logic) is found in Ockham’s Summa Logicae[13][79-80].
3.2 The Logic of the Prior Analytics
@@
The four basic kinds of protasis dealt with in the Analytics can be seen as corresponding to four $(1,1)$-predicate term symbols. Syllogisms correspond very naturally to complex sentences in ESOL of a certain structure:
$\forall A,B,C\phi(A,B)\&\psi(B,C)\rightarrow\rho(A,C)$
@@
where $\phi,\psi$ and $\rho$ range among such $(1,1)$- predicate term symbols (the order of the free variables is not necessarily the order of the free variables in their application to the corresponding symbol). For example, the Barbara syllogism corresponds to:
@@
where $\kappa XY$ is to be read as ’all $X$ is $Y$’.
From a formal point of view this is exactly what Aristotle is expressing and what Aristotle makes use of in the structure of the proofs in the Analytics. Aristotle certainly seems to assume the following reductive axiom (or definition) in the proofs:
@@
Thus $\kappa$ is extensional. But Axiom (A) has to be tweaked as there is no direct evidence that Aristotle accepted $\kappa AA$. To remedy this we introduce an extensional second-order equality predicate
@@
As is well known, Aristotle seems to assume that all terms in the syllogistic have non-empty extension. But this evidently refers not to all terms (such as ’unicorn’) but to those considered fundamental or relevant to scientific discourse (though given a term, the examination of whether this term has instances in patently a part of Aristotle’s scientific methodology). Thanks to ESOL we can state this directly (using the $(1)$-predicate term symbol $\rho$) as an axiom
$\forall A(\rho A\rightarrow\exists xAx)$
In this section we implicitly assume that quantifiers over monadic predicate variables $X$ are subject to the restriction of $\rho X$.
Similarly to $\kappa$ we can define $\mu XY$ to express ’some $X$ is $Y$’. We have the following axiom or definition
@@
which states that $\mu$, like $\kappa$, is extensional. Again we mention the problem that we have to exclude $\mu AA$ and this leads us to variant
@@
We can follow the approach of [16] and introduce a second-order function to represent term-negation (de re negation) which we denote by $A^{c}$. But thanks to our ESOL framework we can now define it directly
@@
We can define $\hat{\kappa}XY\leftrightarrow\kappa XY^{c}$ and $\hat{\mu}XY\leftrightarrow\mu XY^{c}$. Alternatively we could introduce $\hat{\kappa}$ and $\hat{\mu}$ directly with the expected axioms, for instance
@@
It is then straightfotward to derive all the assertoric syllogisms consisdered valid by Aristotle as well as to adapt Aristotle’s own counterexamples to demonstrate the invalidity of those he considered invalid.
We now consider briefly how the modal syllogistic could be treated in the framework of ESOL. To this end we consider an additional second-order monadic function $\square X$. This takes a monadic predicate term $T$ and yields a monadic predicate term $\square T$. We posit the following axioms: